A rectangular animal pen will be built using 200 meters of fencing. If one side is 60 meters, find the area of the pen.
I thought ok two side are 60 m each 60x60 1200
however 120x 120 gives me 14,400
total of 15,800 sq ft

Answer by Alan3354(69443)

A rectangular animal pen will be built using 200 meters of fencing. If one side is 60 meters, find the area of the pen.
------------
Perimeter = 2*60 + 2*L = 200
L = 40 meters
Area = 60*40 = 2400 sq meters